OpenStudy (anonymous):

a=713 b=437 Find the gcd(a,b). Find two integers x,y such that 22=xa+yb

OpenStudy (anonymous):

now the gcd is 23 that is easy to fin but how can I find x,y?

jhonyy9 (jhonyy9):

22=713x+437y 22=23(31x+19y) 22/23 =31x+19y 22=713x+437y - so after this i think that you need checking two cases 1. when x=0 than y= 22/437 2. when y=0 than x= 22/713 - so this is just one my opinion from this all (i am not sure that is right sure -- sorry )

OpenStudy (anonymous):

Has to be integers

jhonyy9 (jhonyy9):

sorry you think that after thses calcules will be possibill getting integers too ?

OpenStudy (anonymous):

Now I have 8*713-13*437=23 I know how to do that. Now I would need to solve 1=ax+by

OpenStudy (anonymous):

I believe it has no solutions but Im not sure

jhonyy9 (jhonyy9):

- so i understand your opinion but the result dont need to be 22 ?

OpenStudy (anonymous):

yes it should be :-)

OpenStudy (anonymous):

I dont know...

jhonyy9 (jhonyy9):

22=713x+437y 22=23*31x+23*19y 31x=a 19y=b a+b=22 a=22-b 31x=22-19y - for what integers x,y can you make this equale ?

OpenStudy (anonymous):

You tell me :-)

OpenStudy (anonymous):

I am more and more confident that it has no solutions

OpenStudy (anonymous):

Hope thats right. I wrote that in my exam

OpenStudy (anonymous):

Euclids algorith only work with coprime numbers so I dont think this should work